# Rejoining fragmented ancient bamboo slips with physics-driven deep learning

### v1

@@ Abstract
Bamboo slips are a crucial medium for recording ancient civilizations in East Asia, and offers invaluable archaeological insights for reconstructing the Silk Road, studying material culture exchanges, and global history. However, many excavated bamboo slips have been fragmented into thousands of irregular pieces, making their rejoining a vital yet challenging step for understanding their content. Here we introduce WisePanda, a physics-driven deep learning framework designed to rejoin fragmented bamboo slips. Based on the physics of fracture and material deterioration, WisePanda automatically generates synthetic training data that captures the physical properties of bamboo fragmentations. This approach enables the training of a matching network without requiring manually paired samples, providing ranked suggestions to facilitate the rejoining process. Compared to the leading curve matching method, WisePanda increases Top-50 matching accuracy from 36% to 52%. Archaeologists using WisePanda have experienced substantial efficiency improvements (approximately 20 times faster) when rejoining fragmented bamboo slips. This research demonstrates that incorporating physical principles into deep learning models can significantly enhance their performance, transforming how archaeologists restore and study fragmented artifacts. WisePanda provides a new paradigm for addressing data scarcity in ancient artifact restoration through physics-driven machine learning.
@@
Figure 1: The challenge of bamboo slip rejoining in archaeological research. The excavated bamboo slip fragments from the Qin dynasty (221–206 BCE) arranged in a spiral pattern (left) illustrate the overwhelming scale of the rejoining challenge faced by archaeologists—with thousands of fragments at single excavation sites, each potentially matching with any other piece. A magnified view of a representative fragment (upper right) reveals the characteristic irregular breakage pattern along the fracture edge, while an example of a reconstructed complete slip (lower right) demonstrates how proper rejoining establishes textual content continuity, which is essential for the interpretation of ancient texts. The complex morphology of the fractures, coupled with deterioration from millennia underground, creates a matching problem of extraordinary complexity that has traditionally required intensive manual effort, with experts sometimes spending weeks to successfully match a single pair of fragments.
Bamboo slips, serving as a fundamental medium for documenting ancient East Asian civilizations, contain invaluable historical records spanning philosophy, law, and social life period [1, 2, 3, 4, 5]. Their durability has enabled these artifacts to survive millennia underground while retaining legible content [6], offering scholars unprecedented insights into historical societies. However, the excavation of these delicate artifacts presents a critical challenge [7] - many bamboo slips have been fragmented into thousands of pieces (Figure 1), significantly complicating efforts to reconstruct and interpret their content. This fragmentation creates a fundamental obstacle in accessing the wealth of historical knowledge contained in these artifacts.
The rejoining of fragmented bamboo slips represents one of the most challenging problems in archaeological preservation and cultural heritage studies [8]. The difficulty stems from multiple factors: First, the enormous number of potential fragment combinations makes manual matching extremely time-consuming - for instance, The Qin bamboo slips from the Shuihudi site comprise fragments numbering in the tens of thousands, with each piece potentially matching any of the others [9]. Second, environmental factors like moisture and pressure have caused extensive physical deterioration [10, 11], distorting the original shapes and surfaces. Third, the unique fiber structure of bamboo creates complex fracture patterns that traditional curve-matching approaches struggle to analyze effectively. These challenges have made the restoration process highly labor-intensive, with experts sometimes requiring weeks to successfully match a single pair of fragments [12].
@@
Here we overcome the challenges of fragmented bamboo slip rejoining through physics-driven deep learning. Our approach integrates principles from the physics of fracture with deep neural networks, enabling automated matching of fragmented pieces without relying on manually labeled training data. Deep learning has been successfully used in restoring and attributing ancient texts [13] empowered by large labeled dataset. While recent years have also seen significant progress in fragmented cultural relic restoration [14] with labeled data - from oracle bone reconstruction [15, 16, 15, 17] to manuscript fragment matching [6] - most existing approaches rely on traditional computer vision techniques and struggle with complex degradation patterns [18]. Some works have attempted to use generative models for cultural heritage restoration [17]. While achieving promising results, these methods are constrained by the scarcity of training data and often fail to capture the intricate physical properties of ancient materials [19, 20].
@@
To address these fundamental challenges, we propose WisePanda, a deep learning framework that leverages physical principles to overcome the data scarcity problem inherent in fragmented bamboo slip rejoining tasks. Training such a system presents a unique paradox: while manual fragment rejoining is prohibitively time-consuming - the very problem we aim to solve - this same process would traditionally be required to generate training data for the model. We resolve this dilemma by resorting to the physics of fracture [21, 22]. By modeling the physical properties of bamboo and the process that govern its degradation [23, 10], we generate extensive synthetic training data that captures the essential characteristics of real paired fragment slips (Figure 2). This physics-driven approach enables us to produce large-scale, realistic training data without requiring manual matching efforts, while ensuring the model learns meaningful patterns based on actual material properties rather than superficial features.
@@
Figure 2: Physics-driven modeling of bamboo slip fracture and deterioration. a, The breakage process of bamboo slips showing how fracture propagates across the bamboo’s fiber structure, with the resulting irregular curve (black line) and the corresponding stress field distribution (blue gradient). b, Detailed stress field model illustrating the mathematical relationship between the fracture angles ($\theta_{i-1}$, $\theta_{i}$) and the stress propagation in the $x$-$y$ coordinate system, where the blue dotted circles represent the stress field emanating from the fracture endpoint, showing how stress radiates outward and concentrates at fiber boundaries. c, Time-sequential simulation of the corrosion process, demonstrating how the original fracture pattern (left) changes through environmental exposure (middle) to produce the final deteriorated edge morphology (right), with protruding areas experiencing accelerated degradation. d, Collection of real bamboo slip fracture curves extracted from archaeological samples, exhibiting diverse breakage patterns that serve as reference for model validation. e, Synthetically generated fracture curves produced by our physics-driven model, displaying morphological characteristics highly similar to the real samples in panel d. This approach enables the generation of extensive paired training data that captures both the physical properties of bamboo fragmentation and the effects of long-term degradation, effectively addressing the data scarcity challenge inherent in ancient artifact restoration.
@@
The generation of training data for bamboo slip rejoining requires a comprehensive understanding of both fracture formation and degradation processes [24]. Our physical model builds on two key observations: First, bamboo slips exhibit distinct “transverse” and “longitudinal” fracture modes due to their unique fiber structure [24], as illustrated in Extended Figure E1, with transverse fractures accounting for approximately 70% of cases. Second, these fractures undergo complex corrosion processes over time, significantly altering their original patterns [25, 10, 2]. Building on these insights, we develop a systematic approach to generate synthetic fragment pairs for transverse fractures, which represent the majority of rejoining challenges.
The physics of fracture in bamboo slips is governed by their distinctive vertical fiber arrangement [26, 27, 28, 29], as detailed in Extended Figure E2. When stress extends horizontally across the slip, it creates characteristic fracture patterns as it propagates through consecutive fiber bundles [30, 27, 29]. We model this process through a probabilistic framework where the fracture angle of each fiber bundle is influenced by the stress field generated at the fracture endpoint of the preceding fiber bundle. This physical relationship can be expressed through a probability density function that determines likely fracture paths based on stress distribution and geometric configurations. The corrosion process is then simulated by analyzing geometric exposure - protruding areas and exposed surfaces are more susceptible to degradation through moisture absorption and microorganism activity [31]. To elegantly calculate the exposure area of each fiber, we ingeniously employ ReLU functions that compare the height differences between adjacent fibers, effectively capturing only the protruding portions that are more vulnerable to environmental deterioration. This computational approach allows us to precisely model the progressive and differential deterioration patterns observed in archaeological samples.
To ensure the generated data accurately reflects real bamboo slip characteristics, we optimize model parameters using a genetic algorithm [32, 33]. The algorithm evaluates parameter sets by comparing the distribution of generated fragments against a reference set of 200 real fragment edges through dimensionality reduction and Silhouette analysis [34]. Our analysis shows high similarity between the synthetic and real data distributions. The resulting pipeline enables us to generate extensive training data that captures both the physical properties of fracture formation and the effects of long-term degradation, effectively addressing the data scarcity challenge in rejoining fragmented bamboo slips.
@@
Figure 3: Framework of WisePanda integrating physics-driven deep learning for bamboo slip rejoining. The workflow diagram illustrates the complete pipeline of our system. The process begins with input bamboo slip fragments (top row), where the target fragment is highlighted in pink. In the physical model stage, we implement two key components: a physics-driven fracture engine that generates initial breakage patterns, and a corrosion simulation modifier that replicates long-term degradation effects. The dimensionality reduction module (center) enables comparison between real and synthetic fragments through statistical visualization techniques, facilitating parameter optimization to ensure realistic fracture simulation. The optimized physical model feeds into the training pair synthesizer, which generates matching fragment pairs for training the matching network. This matching network learns to effectively represent and compare fragment features through a multi-layer architecture with specialized embedding capabilities. During inference (red arrows), the system compares the target fragment against all candidates, producing a ranked list of potential matches (bottom row). The ranking transforms thousands of possible matches into a manageable set, with the system highlighting the most probable candidates (position 5 shown in pink represents a correct match). This approach enables archaeologists to focus their verification efforts on the most promising candidates, significantly accelerating the traditionally time-consuming rejoining process.
@@
WisePanda’s architecture features an interpretable pipeline for fragmented bamboo slip rejoining task (Figure 3), using synthesized data with physics of fracture for model training and providing top-k candidates to assist archaeologists. At its core lies a TripletNet-based deep learning network [35], designed to learn effective feature representation that can distinguish between matching and non-matching fragments. The network processes each input as a 64-dimensional vector that captures the detailed features of the fracture edge curves [36, 37, 38]. By leveraging a triplet loss function [39, 40, 41], the network learns to minimize the distance between matching pairs while maximizing the distance between non-matching ones.
@@
In practical deployment, WisePanda operates by predicting matching probabilities between fragments and offers archaeologists a ranked list of the top potential matches for each fragment. This significantly reduces the search space from thousands of candidates to a manageable set [42, 43, 44], thereby streamlining the rejoining process. The network’s output is a value between 0 and 1, indicating the model’s confidence in a match between the input fragments, where a value closer to 1 signifies a higher likelihood of a correct match. Moreover, this ranking approach effectively mitigates the overwhelming task of manually matching fragments by narrowing down the possibilities to the most probable matches [45].
@@
To bridge the gap between computational methods and archaeological research practices, we have developed an intuitive computer-assisted tool that implements WisePanda’s capabilities in a practical workflow (Extended Figure E5). When working with this tool, archaeologists first select a target fragment from their collection. The system then automatically generates a ranked list of potential matching candidates [43, 42, 46], significantly narrowing down the search space from thousands to dozens of fragments. Each suggested match is accompanied by a confidence score derived from WisePanda’s matching network, allowing archaeologists to prioritize their examination of the most promising candidates.
The tool provides a dedicated workspace where archaeologists can visually manipulate and compare fragments through intuitive operations such as dragging, rotating, and fine-tuning alignments. To facilitate detailed examination, the system offers specialized visualization features like layer swapping and edge enhancement, enabling careful inspection of ink traces and fracture patterns All successful matches are saved and documented in a centralized database, creating a growing repository of verified rejoining that contributes to the broader restoration effort. This computer-assisted tool has significantly accelerated the traditional rejoining workflow [47, 44], ensuring archaeologists maintain full control over the final verification of matches while benefiting from AI-powered suggestions.
@@
Datasets. For evaluating WisePanda’s effectiveness, we compiled several datasets of bamboo and wooden slip fragments. Our primary test set (Bamboo236) consists of 118 pairs of matched bamboo slip fragments, carefully verified and provided by archaeological experts who successfully rejoined these fragments through meticulous manual work. These bamboo slip artifacts are part of an ongoing archaeological research project, with the cultural relics yet to be formally published. To simulate more challenging real-world scenarios, we created an extended dataset (Bamboo1350) by introducing 1114 additional interference fragments, expanding the candidate pool to 1,232 total fragments while maintaining the original test pairs. To assess generalization capabilities across different materials, we also compiled wooden slip datasets. The base wooden slip test set (Wood670) contains 335 expert-verified pairs, while its extended version (Wood3833) includes 3163 additional interference fragments. These diverse datasets enable comprehensive evaluation of fragment matching methods under varying conditions and material types.
@@
Evaluation methodology. The evaluation process involved computing similarity scores between each fragment’s fracture edge and all potential matching candidates from the complementary group. Fragments were pre-classified as upper or lower parts, with matching only performed between these distinct groups. We extracted features from both fracture edges of each fragment, ranked the computed similarity scores, and measured the probability of finding the correct match within the Top-k predictions. We conducted comprehensive experiments comparing our approach with several existing methods, including Manual Random Search, curve matching techniques (using Cosine Similarity and Dynamic Time Warping (DTW)) [48], Fast Matching Method (FMM) [49, 50], and Scale Invariant Signature (SIS) [51]. This comparative framework allows us to assess the relative performance of our physics-driven approach against established techniques in the field of fragment matching.
Results and analysis. The performance variations among different methods reveal insights about the nature of the bamboo slip matching problem. In our original test set (Bamboo118), WisePanda demonstrated superior performance across all metrics, achieving a top-1 accuracy of 12.29%, top-5 accuracy of 35.17%, top-10 accuracy of 52.54%, top-20 accuracy of 69.07%, and a top-50 accuracy of 94.07%. Among traditional methods, SIS achieved the second-best performance with a top-1 accuracy of 10.59% and top-50 accuracy of 72.88%, followed by DTW with a top-50 accuracy of 75.42%. FMM and Cosine Similarity performed less effectively, with top-50 accuracies of 65.25% and 17.80% respectively.
@@
To simulate real-world archaeological scenarios where numerous unrelated fragments exist, we expanded our test dataset by introducing additional interference fragments. As shown in Extended Figure E4, when the candidate pool was increased by a factor of approximately 10, WisePanda maintained a top-50 accuracy of 52.54% compared to the original 94.07%, demonstrating robust performance even with significantly increased search space.
Method Top1(%) Top5(%) Top10(%) Top20(%) Top50(%) Top100(%)
@@
DTW [48] 7.20 22.46 30.93 49.58 75.42 -
FMM [50] 8.90 18.64 26.27 38.14 65.25 -
SIS [52] 10.59 22.03 30.93 47.03 72.88 -
@@
DTW 3.81 12.29 14.41 20.34 32.20 37.71
FMM 6.36 11.44 13.14 16.10 25.00 33.47
SIS 7.20 15.25 19.07 24.15 36.02 48.73
@@
Wood WisePanda[Wood670] 5.38 15.37 26.42 38.36 64.18 86.27
WisePanda[Wood3833] 2.69 9.55 13.88 21.64 34.78 48.36
Table 1: Comparison of fragment matching methods across different test scenarios. Results show Top-k matching accuracy (percentage of correctly identified matches within the top k candidates) across bamboo and wooden slip datasets. The original bamboo test set (Bamboo118) contains 118 paired fragments, while the extended version (Bamboo1114) includes 1,114 fragments with additional interference fragments to create a more challenging matching environment. Similarly, for wooden slips, we compare performance on the original test set (Wood335) and an extended version (Wood3163). As demonstrated in Extended Figure E4, WisePanda consistently outperforms traditional curve matching approaches in both material types and across varying complexity levels, demonstrating the effectiveness of our physics-driven approach in handling diverse archaeological fragments.
@@
The fundamental challenge in generating training data for bamboo slip matching lies in the irreversible nature of the physical degradation process [10, 18]. When archaeologists uncover bamboo slip fragments, they observe only the final degraded edges [7]. The original fracture patterns, which are crucial for accurate matching, are permanently altered by centuries of degradation and cannot be directly inferred from the preserved fragments [8, 53]. This irreversibility poses a significant limitation for conventional data-driven approaches [54, 19, 20, 17], which typically require paired samples of ”before-degradation” and ”after-degradation” states for effective training, data that is inherently unavailable in archaeological contexts. Even generative models like GANs suffer from distribution bias when learning only from successfully matched fragments, failing to incorporate the underlying physical principles that govern fracture formation and degradation.
Our physics-driven model addresses this challenge by modeling the forward process of fragmentation and degradation. We first simulate the initial fracture patterns based on the physics of bamboo fiber structures [24, 55, 56, 23], then apply physical degradation processes to these patterns. This sequential modeling captures both the immediate fracture characteristics and their subsequent transformation through environmental exposure. The physical processes governing bamboo deterioration follow specific geometric principles, where protruding areas and exposed surfaces are more susceptible to degradation [10, 18]. By incorporating these physical constraints, WisePanda generates realistic degradation patterns that closely mirror natural deterioration processes, effectively bridging the gap between the unknown original fracture patterns and the observed degraded edges.

### v2

@@ Abstract
Bamboo slips are a crucial medium for recording ancient civilizations in East Asia, and offers invaluable archaeological insights for reconstructing the Silk Road, studying material culture exchanges, and global history. However, many excavated bamboo slips have been fragmented into thousands of irregular pieces, making their rejoining a vital yet challenging step for understanding their content. Here we introduce WisePanda, a physics-driven deep learning framework designed to rejoin fragmented bamboo slips. Based on the physics of fracture and material deterioration, WisePanda automatically generates synthetic training data that captures the physical properties of bamboo fragmentations. This approach enables the training of a matching network without requiring manually paired samples, providing ranked suggestions to facilitate the rejoining process. Compared to the leading curve matching method, WisePanda increases Top-50 matching accuracy from 36% to 52% among more than one thousand candidate fragments. Archaeologists using WisePanda have experienced substantial efficiency improvements (approximately 20 times faster) when rejoining fragmented bamboo slips. This research demonstrates that incorporating physical principles into deep learning models can significantly enhance their performance, transforming how archaeologists restore and study fragmented artifacts. WisePanda provides a new paradigm for addressing data scarcity in ancient artifact restoration through physics-driven machine learning.
{refsection}
@@
Figure 1: The challenge of bamboo slip rejoining in archaeological research. Some excavated bamboo slip fragments from the Qin dynasty (221–206 BCE) arranged in a spiral pattern (left) illustrate the overwhelming scale of the rejoining challenge faced by archaeologists—with thousands of fragments at single excavation sites, each potentially matching with any other piece. A zoomed view of a representative fragment (upper right) reveals the characteristic irregular breakage pattern along the fracture curve. An example of a reconstructed complete slip (lower right) demonstrates how proper rejoining establishes morphological continuity, which is essential for the interpretation of ancient texts. The complex morphology of the fractures, coupled with deterioration from millennia underground, creates a matching problem of extraordinary complexity that has traditionally required intensive manual effort, with experts sometimes spending weeks to successfully match a single pair of fragments.
Bamboo slips, serving as a fundamental medium for documenting ancient East Asian civilizations, contain invaluable historical records spanning philosophy, law, and social life of the period [allan2015buried, liu2024approaching, xiaogan2003bamboo, ling1990formulaic, zheng2006brief]. Their durability has enabled these artifacts to survive millennia underground while retaining legible content, offering scholars unprecedented insights into historical societies. However, the excavation of these delicate artifacts presents a critical challenge [johnson2019archaeological, liu2024approaching] - many bamboo slips have been fragmented into thousands of pieces (Figure 1), significantly complicating efforts to reconstruct and interpret their content. This fragmentation creates a fundamental obstacle in accessing the wealth of historical knowledge contained in these artifacts.
The rejoining of fragmented bamboo slips represents one of the most challenging problems in archaeological preservation and cultural heritage studies [zhou2017identifying]. The difficulty stems from multiple factors: First, the enormous number of potential fragment combinations makes manual matching extremely time-consuming - for instance, The Qin bamboo slips from the Shuihudi site comprise fragments numbering in the tens of thousands, with each piece potentially matching any of the others [Shuihudi1990QinSlips]. Second, environmental factors like moisture and pressure have caused extensive physical deterioration [cha2014micromorphological], distorting the original shapes and surfaces. Third, the sparse recorded characters often make it difficult to capture textual remnants at the fracture curves created by transverse breakage, rendering text-based matching approaches largely ineffective and necessitating reliance primarily on the morphological patterns of fracture curves for identification. The unique fiber structure of bamboo creates complex fracture patterns that traditional curve-matching approaches struggle to analyze effectively. These challenges have made the restoration process highly labor-intensive, with experts sometimes requiring weeks to successfully match a single pair of fragments.
@@
Here we overcome the challenges of fragmented bamboo slip rejoining through physics-driven deep learning. Our approach integrates principles from the physics of fracture with deep neural networks, enabling automated matching of fragmented pieces without relying on manually labeled training data. Deep learning [lecun2015deep] has been successfully used in restoring and attributing ancient texts [assael2022restoring] empowered by large training datasets. Beyond cultural heritage applications, deep learning has also demonstrated remarkable success across diverse scientific domains [silver2017mastering, senior2020improved, jumper2021highly, dafoe2021cooperative, merchant2023scaling, lu2024multimodal, bodnar2025foundation]. While recent years have also seen significant progress in fragmented cultural relic restoration [bickler2021machine] with labeled data - from oracle bone reconstruction [meng2018recognition, zhang2022data] to manuscript fragment matching [zhu2007globally, zheng2024reunion] - most existing approaches rely on traditional computer vision techniques and struggle with complex degradation patterns [li2014investigation]. Some works have attempted to use generative models for cultural heritage restoration [zhang2022data]. While achieving promising results, these methods are constrained by the scarcity of training data and often fail to capture the intricate physical properties of ancient materials [eslami2020review, hou2018novel].
@@
We propose WisePanda, a deep learning framework that leverages physical principles to overcome the data scarcity problem inherent in fragmented bamboo slip rejoining task. Training WisePanda presents a unique paradox: while manual fragment rejoining is prohibitively time-consuming. The very problem we aim to solve - this same process would traditionally be required to generate training data for the model. We resolve this dilemma by resorting to the physics of fracture [anderson2005fracture]. By modeling the physical properties of bamboo and the process that govern its degradation [tan2011mechanical, cha2014micromorphological], we generate extensive synthetic training data that captures the essential characteristics of real paired fragment slips (Figure 2). This physics-driven approach enables us to produce large-scale, realistic paired training data without requiring manual matching efforts, while ensuring the model learns meaningful patterns based on actual material properties rather than superficial features.
@@
Figure 2: Physics-driven modeling of bamboo slip fracture and deterioration. a, The breakage process of bamboo slips showing how fracture propagates across the bamboo’s fiber structure, with the resulting irregular curve composed of black line segments and the corresponding stress field distribution (blue gradient). b, Detailed stress field model illustrating the mathematical relationship between the fracture angles ($\theta_{i-1}$, $\theta_{i}$) and the stress propagation in the $x$-$y$ coordinate system, where the blue dotted circles represent the stress field emanating from the fracture endpoint, showing how stress radiates outward and concentrates at fiber boundaries. c, Time-sequential simulation of the corrosion process, demonstrating how the original fracture pattern (left) changes through environmental exposure (middle) to produce the final deteriorated curve morphologies (right), with protruding areas experiencing accelerated degradation. d, Collection of real bamboo slip fracture curves extracted from archaeological samples, exhibiting diverse breakage patterns that serve as reference for model validation. e, Synthetically generated fracture curves produced by our physics-driven model, displaying morphological characteristics highly similar to the real samples in panel d. This approach enables the generation of extensive paired training data that captures both the physical properties of bamboo fragmentation and the effects of long-term degradation, effectively addressing the data scarcity challenge inherent in ancient artifact restoration.
@@
The generation of training data for bamboo slip rejoining requires a comprehensive understanding of both fracture formation and degradation processes [behera2025review]. Our physical model builds on two key observations: First, bamboo slips exhibit distinct “transverse” and “longitudinal” fracture modes due to their unique fiber structure [behera2025review], as illustrated in Extended Figure E1, with transverse fractures accounting for approximately 70% of cases. Second, these fractures undergo complex corrosion processes over time, significantly altering their original patterns [chen2022state, cha2014micromorphological, liu2024approaching]. Building on these insights, we develop a systematic approach to generate synthetic fragment pairs for transverse fractures, which represent the majority of rejoining challenges.
The physics of fracture in bamboo slips is governed by their distinctive vertical fiber arrangement [dixon2014structure, ramful2020investigation, chen2019fracture], as detailed in Extended Figure E2. When stress extends horizontally across the slip, it creates characteristic fracture patterns as it propagates through consecutive fiber bundles [chen2019fracture, ramful2022investigation]. We model this process through a probabilistic framework where the fracture angle of each fiber bundle is influenced by the stress field generated at the fracture endpoint of the preceding fiber bundle (Figure 2a and b). This physical relationship can be expressed through a probability density function that determines likely fracture paths based on stress distribution and geometric configurations. The corrosion process is then simulated by analyzing geometric exposure - protruding areas and exposed surfaces are more susceptible to degradation through moisture absorption and microorganism activity [blanchette1994assessment]. To efficiently calculate the exposure area of each fiber bundle, we ingeniously employ ReLU functions that compare the height differences between adjacent fiber bundles, effectively capturing only the protruding portions that are more vulnerable to environmental deterioration (Figure 2c). This computational approach allows us to precisely model the progressive and differential deterioration patterns observed in archaeological samples.
To ensure the generated data accurately reflects real bamboo slip characteristics, we optimize model parameters using a genetic algorithm [Goldberg1989]. The algorithm evaluates parameter sets by comparing the distribution of generated fragments against a reference set of 200 real fragment curves through dimensionality reduction and Silhouette analysis [van2008visualizing]. Our analysis shows high similarity between the synthetic and real data distributions. The resulting pipeline enables us to generate extensive paired training data (Figure 2d and e) that captures both the physical properties of fracture formation and the effects of long-term degradation, effectively addressing the data scarcity challenge in rejoining fragmented bamboo slips.
@@
Figure 3: Framework of WisePanda integrating physics-driven deep learning for bamboo slip rejoining. The workflow diagram illustrates the complete pipeline of our system. The process begins with input bamboo slip fragments (top row), where the target fragment is highlighted in pink. In the physical model stage, we implement two key components: a physics-driven fracture engine that generates initial breakage patterns, and a corrosion simulation modifier that replicates long-term degradation effects. The dimensionality reduction module (center) enables comparison between real and synthetic fragments through statistical visualization techniques, facilitating parameter optimization to ensure realistic fracture simulation. The optimized physical model feeds into the training pair synthesizer, which generates matching fragment pairs for training the matching network. This matching network learns to effectively represent and compare fragment curves through a multi-layer architecture with specialized embedding capabilities. During inference (red arrows), the system compares the target fragment against all candidates, producing a ranked list of potential matches (bottom row). The ranking transforms thousands of possible matches into a manageable set, with the system highlighting the most probable candidates (position 5 shown in pink represents a correct match). This approach enables archaeologists to focus their verification efforts on the most promising candidates, significantly accelerating the traditionally time-consuming rejoining process.
@@
WisePanda’s architecture features an interpretable pipeline for fragmented bamboo slip rejoining task (Figure 3), using synthesized data with physics of fracture for model training and providing Top-k candidates to assist archaeologists. At its core lies a TripletNet-based deep learning network [hoffer2015deep], designed to learn effective feature representation that can distinguish between matching and non-matching fragments. The network processes each input as a 64-dimensional vector that captures the detailed features of the fracture curves. By leveraging a triplet loss function [hoffer2015deep, tian2020hynet], the network learns to minimize the distance between matching pairs while maximizing the distance between non-matching ones.
@@
In practical deployment, WisePanda operates by predicting matching probabilities between fragments and offers archaeologists a ranked list of the top potential matches for each fragment. This significantly reduces the search space from thousands of candidates to a manageable set [funkhouser2011learning, brown2012tools], thereby streamlining the rejoining process. The network’s output is a value between 0 and 1, indicating the match probability between a given pair of fragments, where a value closer to 1 signifies a higher likelihood of a correct match. This ranking approach effectively mitigates the overwhelming task of manually matching fragments by narrowing down the possibilities to the most probable matches.
@@
To bridge the gap between computational methods and archaeological research practices, we have developed an intuitive computer-assisted tool that implements WisePanda’s capabilities in a practical workflow (Extended Figure E5). When working with this tool, archaeologists first select a query fragment from the collection. The system then automatically generates a ranked list of potential matching candidates, significantly narrowing down the search space from thousands to dozens of fragments. Each suggested match is accompanied by a matching score given by WisePanda’s matching network, allowing archaeologists to prioritize their examination of the most promising candidates.
The tool provides a dedicated workspace where archaeologists can visually manipulate and compare fragments through intuitive operations such as dragging, rotating, and fine-tuning alignments. To facilitate detailed examination, the system offers specialized visualization features like layer swapping and curve enhancement, enabling careful inspection of fracture patterns, ink traces, and textures (e.g., fiber patterns). All successful matches are saved and documented in a centralized database, creating a growing repository of verified rejoining that contributes to the broader restoration effort. This computer-assisted tool has significantly accelerated the traditional rejoining workflow [brown2012tools], ensuring archaeologists maintain full control over the final verification of matches while benefiting from AI-powered suggestions.
@@
Datasets. For evaluating WisePanda’s effectiveness, we collected several datasets of bamboo and wooden slip fragments. Our primary test set (Bamboo236) consists of 118 pairs of matched bamboo slip fragments, carefully verified and provided by archaeological experts who successfully rejoined these fragments through meticulous manual work (achieved by three archaeological experts within half a year). These bamboo slip artifacts are part of an ongoing archaeological research project, with the cultural relics yet to be formally published. To simulate more challenging real-world scenarios, we created an extended dataset (Bamboo1350) by introducing 1114 additional interference fragments from the same set, expanding the candidate pool to 1,350 total fragments while maintaining the original matched pairs. To assess generalization capabilities across different materials, we also collected wooden slip datasets. The base wooden slip test set (Wood670) contains 335 expert-verified pairs, while its extended version (Wood3833) includes 3163 additional interference fragments. These diverse datasets enable comprehensive evaluation of fragment matching methods under varying conditions and material types.
@@
Evaluation methodology. The evaluation process involved computing similarity scores between each fragment’s fracture curve and all potential matching candidates from the pool. Fragments were pre-classified as upper or lower parts, with matching only performed between these distinct groups. We extracted features from both fracture curves of each fragment, ranked the computed similarity scores, and measured the probability of finding the correct match within the Top-k predictions. We conducted comprehensive experiments comparing our WisePanda with several classical methods, including Manual Random Search, curve matching techniques (using Dynamic Time Warping (DTW)) [berndt1994using], Fast Matching Method (FMM) [frenkel2003curve], and Scale Invariant Signature (SIS) [calabi1998differential]. This comparative framework allows us to assess the relative performance of our physics-driven approach against established techniques in the field of fragment matching.
@@
Results and analysis. The performance variations among different methods reveal insights about the nature of the fragmented bamboo slip matching problem. In our original test set (Bamboo236), WisePanda demonstrated superior performance across all metrics, achieving a Top-1 accuracy of 12.29%, Top-5 accuracy of 35.17%, Top-10 accuracy of 52.54%, Top-20 accuracy of 69.07%, and Top-50 accuracy of 94.07%. Among traditional methods, SIS [calabi1998differential] achieved the second-best performance with a Top-1 accuracy of 10.59% and Top-50 accuracy of 72.88%, followed by DTW [berndt1994using] with a Top-50 accuracy of 75.42%. FMM [frenkel2003curve] performed less effectively, with Top-50 accuracies of 65.25%. To simulate real-world archaeological scenarios where numerous unrelated fragments exist, we expanded our test dataset by introducing additional interference fragments. As shown in Extended Figure E4, when the candidate pool was increased by a factor of approximately 10, WisePanda maintained a Top-50 accuracy of 52.54% compared to the original 94.07%, demonstrating robust performance even with significantly increased search space.
Table 1: Comparison of fragment matching methods across different test scenarios
Method Top-1(%) Top-5(%) Top-10(%) Top-20(%) Top-50(%) Top-100(%)
@@
DTW [berndt1994using] 7.20 22.46 30.93 49.58 75.42 -
FMM [frenkel2003curve] 8.90 18.64 26.27 38.14 65.25 -
SIS [calabi1998differential] 10.59 22.03 30.93 47.03 72.88 -
@@
DTW [berndt1994using] 3.81 12.29 14.41 20.34 32.20 37.71
FMM [frenkel2003curve] 6.36 11.44 13.14 16.10 25.00 33.47
SIS [calabi1998differential] 7.20 15.25 19.07 24.15 36.02 48.73
@@
Wood DTW [berndt1994using][Wood670] 2.69 6.72 11.04 15.82 23.43 28.51
FMM [frenkel2003curve][Wood670] 2.24 6.72 9.25 13.28 21.94 27.61
SIS [calabi1998differential][Wood670] 2.24 7.91 10.75 15.22 23.43 27.91
WisePanda[Wood670] 1.94 10.00 17.31 25.22 32.99 35.22
WisePanda[Wood3833] 1.19 2.84 5.67 9.25 15.67 19.85
  • Results show Top-k matching accuracy (percentage of correctly identified matches within the top k candidates) across bamboo and wooden slip datasets. The original bamboo test set (Bamboo236) contains 118 paired fragments, while the expanded dataset (Bamboo1350) expands the candidate pool with 1,114 additional interference fragments. Similarly, for wooden slips, we compare performance on the original test set (Wood670) and an expanded dataset (Wood3833). As demonstrated in Extended Figure E4, WisePanda consistently outperforms traditional curve matching approaches in both material types and across varying complexity levels, demonstrating the effectiveness of our physics-driven approach in handling diverse archaeological fragments.
@@
The fundamental challenge in generating training data for fragmented bamboo slip matching lies in the irreversible nature of the physical degradation process [cha2014micromorphological, li2014investigation]. When archaeologists uncover bamboo slip fragments, they observe only the final degraded curves [liu2024approaching] (Figure 2c). The original fracture patterns are permanently altered by centuries of degradation and cannot be directly inferred from the preserved fragments [zhou2017identifying]. This irreversibility poses a significant limitation for conventional data-driven approaches [eslami2020review, hou2018novel, zhang2022data], which typically require paired samples of “before-degradation” and “after-degradation” states for effective training data that is inherently unavailable in archaeological contexts. Even generative models like generative adversarial networks and diffusion models suffer from distribution bias when learning only from real fragments, failing to incorporate the underlying physical principles that govern fracture formation and degradation.
Our physics-driven model addresses this challenge by modeling the forward process of fragmentation and degradation. We first simulate the initial fracture patterns based on the physics of bamboo fiber structures [behera2025review, amada1997fiber, low2006mapping, tan2011mechanical], then apply physical degradation processes to these patterns. This sequential modeling captures both the immediate fracture characteristics and their subsequent transformation through environmental exposure. The physical processes governing bamboo deterioration follow specific geometric principles, where protruding areas and exposed surfaces are more susceptible to degradation [cha2014micromorphological, li2014investigation]. By incorporating these physical constraints, WisePanda generates realistic degradation patterns that closely mirror natural deterioration processes, effectively bridging the gap between the unknown original fracture patterns and the observed degraded edges.